Prove that V1"\cap" V2 is a subspace of W if and only if V1"\subseteq" V2 or V2⊆V1 in which V1 and V2 are subspaces of vector space W.


1
Expert's answer
2021-08-11T15:44:41-0400

Let's prove that V1"\bigcap" V2 is subspace in W in all cases without any additional conditions.

Let x,y"\in" V1"\bigcap" V2 (any two elements in intersection) and "\alpha,\beta\in"R (any two real coefficients).We must prove that a linear combination x*"\alpha" +y*"\beta""\in" V1"\bigcap" V2 too. By definition of intersection we have x,y"\in" V1 and x,y"\in" V2, besides, because V1,V2 are subspaces in W we have x*"\alpha" +y*"\beta""\in" V1 and x*"\alpha" +y*"\beta""\in" V2 and using again definition of intersection we have in result x*"\alpha" +y*"\beta""\in" V1"\bigcap" V2 and this means that the proof is finished.